OpenStudy (anonymous):

2. Two bicycles depart from Miami Beach going in opposite directions. The first bicycle is traveling at 10 miles per hour. The second bicycle travels at 5 miles per hour. How long does it take until the bikes are 45 miles apart?

OpenStudy (anonymous):

is the answer 9 hours

OpenStudy (anonymous):

how do u do it?

OpenStudy (anonymous):

is that the right answer

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

see when the bicycles go in opposite direction their speeds are subtracted, and the distance between them is 45 hence time = distance /speed 45/10-5= 9 hrs

ganeshie8 (ganeshie8):

distance = speed * time say, it took "t" hours for them to be 45 miles apart in t hours, First bicycle going at 10 mph, travels a distance : 10 * t in t hours, Second bicycle going at 5 mph, travels a distance : 5 * t then, both of them are 45 miles apart : 10t+5t = 45 => t = 45/15 = 3 hrs